OpenStudy (prupel85):

which two functions are inverse of each other A: f(x)=x, g(x)=-x B: f(x)=2x, g(x)=-1/2x C: f(x)=4x, g(x)= 1/4x D: f(x)=-8x, g(x)=8x

OpenStudy (lord_box):

Remember, the slopes of inverse functions are opposite reciprocals of each other

OpenStudy (prupel85):

so its A

OpenStudy (lord_box):

x is a reciprocal of 1/x. -x is the opposite of x. -1/x is the opposite reciprocal of x.

OpenStudy (prupel85):

ohh so its B

OpenStudy (lord_box):

Yep

TheSmartOne (thesmartone):

f(x)=2x y = 2x x = 2y y = x/2 = (1/2)x can't be B :P